# Large Language Model Unlearning

### v1

@@ Abstract
We study how to perform unlearning, i.e. forgetting undesirable (mis)behaviors, on large language models (LLMs). We show at least three scenarios of aligning LLMs with human preferences can benefit from unlearning: (1) removing harmful responses, (2) erasing copyright-protected content as requested, and (3) eliminating hallucinations. Unlearning, as an alignment technique, has three advantages. (1) It only requires negative (e.g. harmful) examples, which are much easier and cheaper to collect (e.g. via red teaming or user reporting) than positive (e.g. helpful and often human-written) examples required in RLHF (RL from human feedback (Ouyang et al., 2022; Bai et al., 2022)). (2) It is computationally efficient. (3) It is especially effective when we know which training samples cause the misbehavior. To the best of our knowledge, our work is among the first to explore LLM unlearning. We are also among the first to formulate the settings, goals, and evaluations in LLM unlearning. We show that if practitioners only have limited resources, and therefore the priority is to stop generating undesirable outputs rather than to try to generate desirable outputs, unlearning is particularly appealing. Despite only having negative samples, our ablation study shows that unlearning can still achieve better alignment performance than RLHF with just 2% of its computational time.
@@ 1 Introduction
Making sure large language models (LLMs) generate safe outputs that align with human values and policy regulation is currently a major task for LLM practitioners.
The common tasks include the following. (1) Removing Harmful Responses: Because LLMs are trained on the Internet data which contain countless harmful text, LLMs are easy to learn problematic responses. For example, (Zhuo et al., 2023; Bai et al., 2022; Liu et al., 2023) have shown that LLMs can memorize harmful concepts; such responses can cause great harm to users. (2) Erasing Copyrighted Content: We have seen a large number of recent efforts to protect the right of content creators from being used by LLMs (Carlini et al., 2021; Wahle et al., 2022; Lee et al., 2023; Copilot, 2023; Liu et al., 2023). They show LLMs can memorize and leak copyright-protected content. Removing such behaviors learned by the LLMs as requested by the authors is important but is prohibitively expensive if we need to retrain LLMs from scratch. (3) Reducing Hallucinations: LLMs often give factually wrong responses that mislead users. Reducing hallucinations, especially in high-stakes applications, is the key to earning user trust. (4) Protecting User Privacy: Users might choose not to give consent to the LLM service providers for using their data. For example, OpenAI has updated its policy regarding how it would use user data for training models (OpenAI, 2023). Therefore, when users stop giving consent, LLM practitioners need a way of removing the user data from the trained LLMs. (5) Enforcing Policy Compliance: Local community compliance policy can iterate frequently (TikTok, 2023; Twitter, 2023; Facebook, 2023). Practitioners need techniques to quickly remove historical training data that leads to outputs that are no longer policy-compliant.
Though those tasks seem different, the central technical question is identical: How to quickly remove the impact of certain training samples on LLMs? To this end, we study how to perform large language model unlearning. If an LLM learns unwanted (mis)behaviors in its pretraining stage, we aim to unlearn them with samples that represent those problematic behaviors, i.e. with only negative samples.
@@
We summarize the benefits of LLM unlearning. (1) It only requires negative examples that we want the LLM to forget, which are cheaper and easier to collect through user reporting or red teaming than positive examples (that are required in the standard RLHF). In addition, discovering negative examples is highly automatable given the pretrained (unaligned) LLM. (2) It is computationally efficient; the cost is similar to finetuning LLMs. (3) Unlearning is particularly efficient in removing unwanted behaviors if practitioners already know which training samples cause them. Given the specific negative samples, it is more effective to remove their impact directly than to do so indirectly by leveraging positive samples (e.g. in RLHF) – if the goal is to not generate undesirable outputs, e.g. generating non-harmful outputs111e.g. nonsensical strings or responses unrelated to prompts. rather than helpful outputs. If we only have limited resources, unlearning provides a promising alternative to RLHF to align LLMs when the first priority is to stop LLMs from generating undesirable outputs since undesirable outputs often cause far more damage than what can be offset by the benefits of desirable outputs.
@@
In this work, we show three successful examples of LLM unlearning: (1) After the LLM learns harmful behaviors from its training data, we want it to stop generating harmful responses (Figure 1). It is similar to the conventional RLHF scenario except the goal is to generate non-harmful responses rather than helpful responses because it is the best we can expect when given only negative samples. (2) After the LLM is trained on copyright-protected content, and the author requests practitioners to remove it, we want to do so without retraining the LLM from scratch (which is forbiddenly costly). (3) If the LLM learns wrong facts in its training data, i.e. “hallucinations,” we want the LLM to forget them.
@@
Unlearning LLMs is different from the traditional unlearning on classification models, and it is more challenging for several reasons. (1) An LLM’s output space is much larger than the label class in classification, and its possible outcomes vastly outnumber the classification. In classification, the definition of unlearning is defined in a more clear-cut way: as long as samples are classified into (or not into) certain classes. However, behaviors are much more ill-defined when the outputs are natural language rather than predicted labels. (2) Given the size of LLMs, the efficiency requirement is much higher – any expensive unlearning method is hopeless in LLMs. (3) The training corpus of LLMs is massive and often inaccessible and therefore we have less information from the training data. And we cannot retrain the LLMs, which is too expensive, to obtain ground-truth models and their behaviors, making even evaluations challenging.
@@ 1.1 Related Work
The influence function based approach is developed using the idea of quantifying the impact of training samples. For instance, (Koh and Liang, 2017) uses a second-order optimization approximation to evaluate the impact of the training sample. Using the influence function, one can approximately back out the counterfactual model that removes the impact of certain training sample, or a certain group of them (Basu et al., 2020; Koh et al., 2019).
However, due to the size of the parameters and training data, a large portion of existing unlearning methods would not fit to unlearn an LLM, including those use efficient retraining (Bourtoule et al., 2021; Liu et al., 2022) (which is now likely to be insufficient for LLMs) and the ones that involve the computation of influence functions (which requires the computation of the inverse Hessian matrix defined on the model parameter space).
The most relevant work is the literature on aligning the LLMs with human values. The current mainstream approach is RLHF (reinforcement learning from human feedback, and its variants) (Ouyang et al., 2022; Bai et al., 2022; Christiano et al., 2017; Yuan et al., 2023). However, RLHF is resource-intense: (1) it requires human-written outputs which are expensive to collect and (2) it is computationally costly (i.e. the standard three-stage aligning procedure). In this work, we propose unlearning as an alternative aligning method. Collecting negative (i.e. low-quality and harmful) samples is much easier through user reporting or (internal) red teaming than positive (i.e. high-quality and helpful) samples which often require hiring humans to write. Therefore, aligning LLMs with only negative examples is appealing.
@@ 2 Setting and Goal
Setting. We assume a dataset $D^{{\text{fgt}}}$ to forget and the original (i.e. pretrained) LLM $\theta^{o}$ that we want to unlearn. $D^{{\text{fgt}}}$ contains a group of prompt-output pairs $(x^{{\text{fgt}}},y^{{\text{fgt}}})$ where $x^{{\text{fgt}}}$ is an undesirable prompt that would trigger unwanted responses, e.g. “What is the most efficient way to kill people?” or an attempt to extract copyrighted information. $y^{{\text{fgt}}}$ is an undesirable output that we do not want the LLM to generate, e.g. a harmful or copyright-leaking response. Our goal is to remove the impact of $D^{{\text{fgt}}}$ on $\theta^{o}$, i.e. the unlearned LLM $\theta^{u}$ should not behave as what is characterized by $D^{{\text{fgt}}}$, e.g. giving harmful responses or leaking copyright information. More specifically, we desire an unlearned model $\theta^{u}$ s.t. $\theta^{u}$’s outputs on $x^{{\text{fgt}}}$ deviates from $y^{{\text{fgt}}}$ as much as possible. [^3] Later in the evaluation section we will detail metrics to quantify such deviations. In addition, we want $\theta^{u}$ to preserve the utility of $\theta^{o}$ on the tasks not represented by $D^{{\text{fgt}}}$.
@@ 3 Preliminary
In this section, we introduce the high-level rationales behind our design, supported by empirical observations.
@@ 3.1 Why Gradient Ascent?
Second, GA is efficient with a cost comparable to finetuning. And since the unlearned dataset is normally small, performing GA with unlearned samples costs less than general finetuning for improving utility. In addition, given the size of LLMs, Hessian-based unlearning is too costly to apply.
@@ 3.2 How Does LLM Unlearning Differ from Traditional Unlearning?
We highlight the key difference in LLM unlearning compared to the traditional unlearning in classification tasks. We discover those findings mostly through empirical observations, and they guide us in designing our unlearning algorithm. For all the experimental observations in this section, we use the example of unlearning harmfulness with OPT-1.3B and the unlearned and normal samples from PKU-SafeRLHF (Ji et al., 2023) and TruthfulQA respectively (Lin et al., 2021).
@@
[…]
@@
[…]
@@
[…]
@@ 4 Method
Let $h_{\theta}(x,y_{<i}):=\mathbb{P}(y_{i}|(x,y_{<i});\theta)$ be the predicted probability of the token $y_{i}$ by an LLM $\theta$ conditioned on the prompt $x$ and the already generated tokens $y_{<i}:=[y_{1},...,y_{i-1}]$.555if $i=1$, then $y_{<i}$ is the empty sequence. Given a prompt-output pair $(x,y)$ and the LLM $\theta$, the loss on $y$ is:
@@
$\mathcal{L}_{{\text{fgt}}}$ $:={\color[rgb]{1,0,0}-}\sum_{(x^{{\text{fgt}}},y^{{\text{fgt}}})\in D^{{\text{fgt}}}}{L(x^{{\text{fgt}}},y^{{\text{fgt}}};\theta_{t})}$ (4)
@@
[…]
We highlight two designs in our method. (1) We find that performing gradient ascent or decent on the output (i.e. $y$) part only is much more effective than on both prompt and output (i.e. $(x,y)$). In other words, the loss should be only computed on the tokens in $y$ conditioned on $x$, excluding the tokens in $x$, i.e. Eqn(3). (2) Adding $\mathcal{L}_{{\text{rdn}}}$ has two advantages. First, it helps the LLM forget the learned undesirable outputs on $x^{{\text{fgt}}}$ by forcing it to predict random outputs. Second, we find empirically that it helps us preserve the normal utility (Section 5).
[…]
5 Experiments
We include three applications of unlearning: (1) Unlearning the harmfulness of outputs responding to harmful prompts, (2) Unlearning copyright-protected contents requested by creators after they have been trained into LLMs, and (3) Unlearning hallucinated outputs. In addition, we also compare our method to RLHF.
5.1 Evaluation Design
We start with formulating our evaluation design. In general, the evaluation of unlearning LLM outputs is more challenging than classification models because LLMs’ output space is much larger and less well-defined. Broadly speaking, our evaluation metrics fall into two categories: (1) performance on the unlearned samples and (2) utility on the remaining samples. For the first consideration, since we want the effectiveness of unlearning to generalize to unseen samples rather than just unlearned samples, we need to test both unlearned and unseen prompts that would cause misbehaviors. We discuss both categories in the following.
Unlearning Performance: We measure the following metrics on the outputs generated given both unlearned prompts that cause unwanted misbehaviors on LLMs as well as unseen prompts that are similar to the exactly unlearned prompts. Note that unlearned prompts might or might not exactly exist in the LLM’s training data. For example, if we want to unlearn a concept, e.g., harmfulness, then the unlearned prompts (and the undesirable outputs) do not need to exactly belong to the LLM’s training data. On the other hand, if we want to unlearn the previously learned copyrighted data, then the unlearned samples often belong to the training set.

### v2

@@ Abstract
We study how to perform unlearning, i.e. forgetting undesirable misbehaviors, on large language models (LLMs). We show at least three scenarios of aligning LLMs with human preferences can benefit from unlearning: (1) removing harmful responses, (2) erasing copyright-protected content, and (3) reducing hallucinations. Unlearning, as an alignment technique, has three advantages. (1) It only requires negative (e.g. harmful) examples, which are much easier and cheaper to collect (e.g. via red teaming or user reporting) than positive (e.g. helpful and often human-written) examples required in RLHF (reinforcement learning from human feedback). (2) It is computationally efficient; the cost is comparable to light supervised finetuning. (3) It is especially effective when we know which training samples cause the misbehavior. To the best of our knowledge, our work is among the first to explore LLM unlearning. We are also among the first to formulate the settings, goals, and evaluations in LLM unlearning. We show that if practitioners only have limited resources, and therefore the priority is to stop generating undesirable outputs rather than to try to generate desirable outputs, unlearning is particularly appealing. Despite only having negative samples, our ablation study shows that unlearning can still achieve better alignment performance than RLHF with just 2% of its computational time.
@@ 1 Introduction
Making sure large language models (LLMs) generate safe outputs that align with human values and policy regulation is currently a major task for LLM practitioners. The common tasks include the following:
  1. 1.
    Removing Harmful Responses: Since LLMs are trained on the Internet data which contain countless harmful text, they are easy to learn problematic responses. For example, (Zhuo et al., 2023; Bai et al., 2022; Liu et al., 2023) have shown that LLMs can memorize harmful concepts; such responses can cause great harm to users.
  2. 2.
    Erasing Copyrighted Contents: The tension between data owners (e.g., authors) and LLM service providers is escalating, leading to legislation such as legal disputes involving OpenAI, Meta, and New York Times (Small, 2023; Grynbaum and Mac, 2023; Copilot, 2023). We have also seen a large number of recent works that show LLMs can memorize and leak copyright-protected information (Carlini et al., 2021; Wahle et al., 2022; Lee et al., 2023; Liu et al., 2023). Removing such behaviors learned by the LLMs as requested by the authors is important but is prohibitively expensive if we need to retrain LLMs from scratch.
  3. 3.
    Reducing Hallucinations: LLMs often give factually wrong responses that mislead users. Reducing hallucinations, especially in high-stakes applications, is the key to earning user trust.
  4. 4.
    Protecting User Privacy: Users might stop giving consent to the LLM service providers for using their data. When it happens, LLM practitioners need a way of removing the old user data from the trained LLMs.
  5. 5.
    Enforcing Policy Compliance: Local community compliance policy can iterate frequently (TikTok, 2023; Twitter, 2023; Facebook, 2023). Practitioners need techniques to quickly remove historical training data that leads to outputs that are no longer policy-compliant.
Though those tasks seem different, the central technical question is identical: How to quickly remove the impact of training samples on LLMs? To this end, we study how to perform large language model unlearning. If an LLM learns unwanted misbehaviors in its pretraining stage, our goal is to unlearn them with samples that represent those problematic behaviors, i.e. with only negative samples.
@@
We summarize the benefits of LLM unlearning. (1) It only requires negative examples that we want the LLM to forget, which are cheaper and easier to collect through user reporting or red teaming than positive examples, which are required in the standard RLHF. In addition, discovering negative examples is highly automatable given the pretrained (i.e. unaligned) LLM. (2) It is computationally efficient; the cost is similar to finetuning LLMs. (3) Unlearning is particularly effective in removing unwanted behaviors if practitioners already know which training samples cause them. Given the specific negative samples, it is more efficient to remove their undesirable impact directly than to do so indirectly by relying on positive samples (e.g. in RLHF) – if the goal is to stop generating undesirable outputs, e.g. generating non-harmful outputs, as opposed to generating helpful outputs, as is the case in RLHF.
We elaborate on the last benefit, which relates to our scenario. We argue that if practitioners only have limited resources, meaning (1) they do not have the budget to hire humans to write desirable outputs (as required in RLHF) and (2) they have limited computational resources, then the first priority should be stopping LLMs from generating undesirable (e.g. harmful) outputs rather than trying to make them generate desirable outputs (e.g. standard polite responses like “As an AI language model…”). This is because undesirable outputs often cause far more damage than what can be offset by the benefits of desirable outputs. If a user asks an LLM hundreds of questions, and he or she gets only one undesirable answer, the trust might be lost immediately – no matter how many desirable responses the LLM could have given them later. It takes years to build trust and only seconds to destroy. We, therefore, argue generating non-harmful outputs, e.g. nonsensical strings or responses unrelated to prompts, has a higher priority in low-resource alignment scenarios.
@@
In this work, we show three successful examples of LLM unlearning. (1) After the LLM learns harmful behaviors from its training data, we want it to stop generating harmful responses (Figure 1). It is similar to the conventional RLHF scenario except the goal is to generate non-harmful responses rather than helpful responses because it is the best we can expect when given only negative samples. (2) After the LLM is trained on copyright-protected content, and the author requests practitioners to remove it, we want to do so without retraining the LLM from scratch (which is forbiddenly costly). (3) If the LLM learns wrong facts in its training data, i.e. “hallucination,” we want the LLM to forget them.
@@
Unlearning LLMs is different from the traditional unlearning on classification models, and it is more challenging for several reasons. (1) An LLM’s output space is much larger than the label class in classification, and its possible outcomes vastly outnumber the classification. In classification, the definition of unlearning is defined in a more clear-cut way: as long as samples are classified into (or not into) certain classes. However, behaviors are much more ill-defined when the outputs are natural language rather than predicted labels. (2) Given the size of LLMs, the efficiency requirement is much higher – any expensive unlearning method is hopeless in LLMs. (3) The training corpus of LLMs is massive and often inaccessible and therefore we have less information from the training data. In addition, we cannot retrain the LLMs, which is too expensive, to obtain ground-truth models and their behaviors, making even evaluations challenging.
@@ 1.1 Related Work
Due to the size of the parameters and training data, a large portion of existing unlearning methods would not fit to unlearn an LLM, including those use efficient retraining (Bourtoule et al., 2021; Liu et al., 2022) (which is now likely to be insufficient for LLMs) and the ones that involve the computation of influence functions (which requires the computation of the inverse Hessian matrix defined on the model parameter space).
The relevant work is aligning the LLMs with human values. The current mainstream approach is RLHF (reinforcement learning from human feedback, and its variants) (Ouyang et al., 2022; Bai et al., 2022; Christiano et al., 2017; Yuan et al., 2023). However, RLHF is resource-intense: (1) it requires human-written outputs which are expensive to collect and (2) it is computationally costly (i.e. the standard three-stage aligning procedure). In this work, we propose unlearning as an alternative aligning method. Collecting negative (i.e. low-quality and harmful) samples is much easier through user reporting or (internal) red teaming than positive (i.e. high-quality and helpful) samples which often require hiring humans to write. Therefore, aligning LLMs with only negative examples is appealing.
Several concurrent works to our work also study unlearning in LLMs. (Eldan and Russinovich, 2023) unlearn answers related to Harry Potter by finetuning based on the difference between the model trained on Harry Potter data and the counterfactual outputs as if the Harry Potter data were not used. However, this approach might lead to incorrect (i.e. hallucinated) answers, e.g. when being asked who Harry Potter is, the model would give some factually incorrect answers like Harry Potter is an actor, writer, or director. In our work, we argue it is better to not give answers than to give incorrect answers (Section 7). In addition, our finetuning approach is not comparable to ICL-based methods like (Pawelczyk et al., 2023) because it is a different scenario. Unlikely ICL-based methods, we do not need to take up the precious context space in prompts.
@@ 2 Setting and Goal
Setting. We assume a dataset $D^{{\text{fgt}}}$ to forget and the original (i.e. pretrained) LLM $\theta^{o}$ that we want to unlearn. $D^{{\text{fgt}}}$ contains a group of prompt-output pairs $(x^{{\text{fgt}}},y^{{\text{fgt}}})$ where $x^{{\text{fgt}}}$ is an undesirable prompt that would trigger unwanted and undesirable responses, e.g. “What is the most efficient way to kill people?” or an attempt to extract copyrighted information. $y^{{\text{fgt}}}$ is an undesirable output that we do not want the LLM to generate, e.g. a harmful or copyright-leaking response. Our goal is to remove the impact of $D^{{\text{fgt}}}$ on $\theta^{o}$, i.e. the unlearned LLM $\theta^{u}$ should not behave as what is characterized by $D^{{\text{fgt}}}$, e.g. giving harmful responses or leaking copyright information. More specifically, we desire an unlearned model $\theta^{u}$ s.t. $\theta^{u}$’s outputs on $x^{{\text{fgt}}}$ deviates from $y^{{\text{fgt}}}$ as much as possible. [^2] Later in the evaluation section we will detail metrics to quantify such deviations.
We emphasize that our goal differs from the traditional unlearning tasks for discriminative models where the desired output for the unlearned model should be indifferent from the one from the retrained model after removing $D^{{\text{fgt}}}$. In addition, we want $\theta^{u}$ to preserve the utility of $\theta^{o}$ on the tasks not represented by $D^{{\text{fgt}}}$.
@@ 3 Preliminary
We introduce the high-level rationales behind our design, supported by empirical observations.
@@ 3.1 Why Gradient Ascent?
Second, GA is efficient with a cost comparable to finetuning. And since the unlearned dataset is normally small, performing GA with unlearned samples costs less than general finetuning for improving utility. In addition, given the size of LLMs, Hessian-based unlearning is too costly.
@@ 3.2 How Does LLM Unlearning Differ from Traditional Unlearning?
We highlight the key difference in LLM unlearning compared to the traditional unlearning in classification tasks. We discover those findings mostly through empirical observations, and they guide us in designing our unlearning algorithm. For all the experimental observations in this subsection, we use the example of unlearning harmfulness with OPT-1.3B and the unlearned and normal samples from PKU-SafeRLHF (Ji et al., 2023) and TruthfulQA respectively (Lin et al., 2021).
@@
Key Difference 1: Both training and validation loss on the unlearned samples have limited indications of unlearning effectiveness. For example, when we apply gradient ascent (GA), even when the loss on the unlearned samples rises to as high as $60+$ after unlearning for $\sim 200$ batches (Figure 2), the LLM still outputs harmful responses to harmful prompts (Table 1). This is not observed in traditional unlearning, where the losses on the forget samples are often strong indicators of the unlearning performance.
@@
Solution 1: We find continuing to unlearn after the loss on harmful samples rises dramatically is necessary for unlearning effectiveness. For example, although the loss on harmful samples already looks promising after unlearning $\sim 200$ batches, we find the LLM only stops outputting harmful responses after $\sim 1000$ batches (Table 1). We also propose an additional loss that randomly mismatches between $x^{{\text{fgt}}}$ and its response to facilitate the forgetting of $y^{{\text{fgt}}}$ (See Section 4).
Key Difference 2: Performance on normal prompts deteriorates easily after unlearning. We find that preserving performance on normal samples is generally harder to achieve than forgetting. For example, with GA, it is often not hard to make the LLM output random responses. [^3] Mostly random characters, e.g. special characters, such as whitespace, new line, non-ASCII characters, or foreign languages. However, the LLM is likely to also generate nonsensical outputs on normal response. Table 2 shows the example of nonsensical outputs after unlearning with gradient ascent on $\sim 1000$ batches. Although the LLM stops generating harmful responses on the harmful prompts, it also generates nonsensical outputs on normal prompts, destroying the LLM’s utility.
@@
Solution 2: We empirically find that merely optimizing the cross-entropy loss on a normal dataset does not maintain the normal performance well. Like existing work in RLHF (Ouyang et al., 2022; Touvron et al., 2023; Zheng et al., 2023; Holtzman et al., 2019), we find that minimizing the divergence between the output on $x^{{\text{nor}}}$ from the unlearned LLM and the original LLM works the best. (See Section 4.)
Key Difference 3: The format (e.g. Q&A, book text, chat, multiple choice etc.) of $D^{{\text{nor}}}$ (for guiding the LLMs to preserve utility on normal tasks) has a large impact on the normal performance. When the format of $D^{\text{nor}}$ and $D^{\text{fgt}}$ differ substantially, the unlearned LLM can learn a shortcut that decides what to output by the format of the prompt only, and therefore does not truly unlearn the concept.
Solution 3: To maintain the normal performance, we find that choosing the format of $D^{{\text{nor}}}$ to be the same with $D^{{\text{fgt}}}$ (e.g. if $D^{{\text{fgt}}}$ is Q&A, then $D^{{\text{nor}}}$ should also be Q&A) can better help preserve normal utility.
@@ 4 Method
Let $h_{\theta}(x,y_{<i}):=\mathbb{P}(y_{i}|(x,y_{<i});\theta)$ be the predicted probability of the token $y_{i}$ by an LLM $\theta$ conditioned on the prompt $x$ and the already generated tokens $y_{<i}:=[y_{1},...,y_{i-1}]$.444if $i=1$, then $y_{<i}$ is the empty sequence. For a prompt-output pair $(x,y)$ and LLM $\theta$, the loss on $y$ is:
@@
$\mathcal{L}_{{\text{fgt}}}:=-\sum_{(x^{{\text{fgt}}},y^{{\text{fgt}}})\in D^{{\text{fgt}}}}{L(x^{{\text{fgt}}},y^{{\text{fgt}}};\theta_{t})}$ (4)
@@
We explain each loss. Eqn(4) is the gradient ascent (GA) loss to forget the unlearned samples. Note we compute it on $y_{{\text{fgt}}}$ only, as indicated in Eqn(3). Eqn(5) forces the LLM to predict a random output $y^{{\text{rdn}}}$ on the unlearned $x^{{\text{rdn}}}$. This term reinforces the forgetting of prompt $x^{{\text{fgt}}}$ by adding irrelevance into the predicted outcome, with the similar insight of label smoothing (Müller et al., 2019) in classification. Eqn(6) is to preserve the normal utility by comparing it with the original LLM (Key Difference 2). Note that we use forward KL (which is typically used in supervised learning) instead of reverse KL (which is typically used in sampling, e.g. RLHF) because it forces the distribution of the unlearned model to cover all the areas of space of the original LLM (Murphy, 2022).
We highlight two designs in our method. (1) We find that performing gradient ascent or decent on the output (i.e. $y$) part only is much more effective than on both prompt and output (i.e. $(x,y)$). In other words, the loss should be only computed on the tokens in $y$ conditioned on $x$, excluding the tokens in $x$, i.e. Eqn(3). (2) Adding $\mathcal{L}_{{\text{rdn}}}$ has several advantages. First, it helps the LLM forget the learned undesirable outputs on $x^{{\text{fgt}}}$ by forcing it to predict random outputs. Second, we find empirically that it helps us preserve the normal utility (Section 6). Last, it can stabilize the unlearning performance when the gradient on $(x,y)$ is small. We include the detailed analysis in Appendix A.
We incorporate the three key lessons in Section 3.2. (1) We continue to unlearn after we have observed the loss on forgetting samples raises to an abnormally high level, for 3x-5x more batches. (2) To preserve normal utility, we minimize the KL divergence on predicted distribution on $x^{{\text{fgt}}}$ between the original and the unlearned LLM, i.e. Eqn(6). (3) We choose $D^{{\text{nor}}}$ to be the same format as $D^{{\text{fgt}}}$, e.g. to unlearn the harmful data from PKU-SafeRLHF which is in the format of Q&A, we use TruthfulQA as the normal data.
5 Evaluation Design
Broadly speaking, our evaluation metrics fall into two categories: (1) performance on the unlearned samples and (2) utility on the remaining samples.
Unlearning Performance: Since we want the effectiveness of unlearning to generalize to unseen samples rather than just unlearned samples, we need to test both unlearned and unseen prompts that would cause misbehavior. We measure the following metrics on the outputs generated given both unlearned prompts that cause unwanted misbehaviors on LLMs as well as unseen prompts that are similar to the exactly unlearned prompts: [^5] Note that unlearned prompts might or might not exactly exist in the LLM’s training data. For example, if we want to unlearn a concept, e.g., harmfulness, then the unlearned prompts (and the undesirable outputs) do not need to exactly belong to the LLM’s training data. On the other hand, if we want to unlearn the previously learned copyrighted data, then the unlearned samples often belong to the training set.